# Conformal prediction for causal effects of continuous treatments

### v2

@@ Appendix B Algorithm
In our main paper, we presented an algorithm for computing the prediction intervals if the propensity score is unknown. Below, we state a second algorithm that is applicable if the propensity score is known. In this case, a convex optimization solver can be used.
@@ 21      $S^{\ast}\leftarrow\frac{S_{\mathrm{up}}+S_{\mathrm{low}}}{2}$;
Computational complexity: The complexity of running our algorithms depends heavily on the employed optimization solver with complexity $\sigma_{s}(n_{c})$ (e.g., polynomial complexity for suitable convex solvers) and the size of the calibration dataset $n_{c}$. This might become costly for large-scale calibration datasets in practice. The outer algorithm has a time complexity of at most $O(\log(\frac{S_{up}-S_{low}}{\varepsilon})+1)$. Overall, our algorithm has a fixed complexity of $O(\log(\frac{S_{up}-S_{low}}{\varepsilon})+\sigma_{s}(n_{c}))$. The complexity of deriving intervals through MC-dropout or ensemble methods depends, however, on the number of MC samples or models, respectively. The latter thus scales with the precision of the intervals, which might be difficult to control. In our work, we use optimization as a tool to provide conformal prediction intervals. Future research should focus on developing more efficient optimization algorithms for this task.
Appendix C Proofs
C.1 Proofs of the supporting lemmas
In the following, we prove Lemma 2 and Lemma 3 from our main paper.
Proof of Lemma 2
@@ C.1 Proofs of the supporting lemmas
Proof of Lemma 3
We first prove that the problem equation PS fulfills the linear independence constraint qualifications. For all $i=m+1,\ldots,n+1$, we denote the constraints of problem equation PS as
@@
As a result, the constraints are linearly independent. This property suffices for the KKT conditions to hold at any (local) optimum of equation PS.To furthermore show that the KKT conditions are also sufficient for a global optimum, we show that equation PS is Type-I invex. An optimization problem with objective function $f(x)$ and constraints $g(x)<=0$ with $x\in\mathbb{R}^{n+1}$ is Type-I invex at $x_{0}$, if there exists $\nu(x,x_{0})\in\mathbb{R}^{n+1}$, such that
@@
(Hanson & Mond, 1987). In problem equation PS, the gradients of the objective function and of each constraint $h_{i}$ for all $i,j=m+1,\ldots,n+1$ at $x_{0}$ are given by
@@
the definition of Type-I invexity holds for equation PS. Thus, the KKT conditions are also sufficient for a global optimum.
C.2 Proof of Theorem 1
We prove Theorem 1 in three steps: (i) We show that function class :=$\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$ indeed satisfies Eq. equation 2 for the intervention $a^{\ast}=a+\Delta_{A}$ and rewrite Eq. equation 5 as a convex optimization problem. (ii) We retrieve the corresponding dual problem, derive a dual prediction set, and show the equality of the coverage guarantee of the dual and the primal prediction sets. (iii) We derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. For further theoretical background on the idea of the proof, we refer to Gibbs et al. (2023).
@@ C.2 Proof of Theorem 1
Observe that $\mathbb{E}[f(A,X)]=\theta$ for all $f\in\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. Therefore, Eq. equation 2 simplifies to
@@
Thus, $\mathcal{F}$ satisfies the propensity shift from Eq. equation 2 for the soft intervention $a^{\ast}=a+\Delta_{A}$. Following Lemma 1, we thus aim to find
@@
For a reference, see (Gibbs et al., 2023). The Lagrangian of the primal problem states
@@
Recall from Lemma 1 that, for
@@
With the definition of $S^{\ast}$ as the maximum optimizer $\eta_{n+1}^{S}$ with $\eta_{n+1}^{S}<1-\alpha$ and Lemma 5, it follows that
@@
C.3 Proof of Theorem 2
We follow the same outline as in the proof of Theorem 1 in Section C.2. In Lemma 2, we motivated the functional class of distribution shifts. Therefore, it is left to prove the coverage guarantee of $C_{n+1}$.
@@ C.3 Proof of Theorem 2
Coverage guarantees. As in C.2, we fix some $\sigma>0$ and $c_{a}\in[\frac{1}{M},M]$ to obtain a specific $f(a,x):=\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp{\left(-\frac{(a_{i}-a^{\ast})^{2}}{2\sigma^{2}}\right)}}{\hat{\pi}(a_{i}\mid x_{i})}$. We further denote $\hat{g}(a,x)\in\mathcal{F}$ the optimal solution given by the optimal values $\hat{\sigma}$ and $\hat{c_{a}}$.
With the definition of $S^{\ast}$ as the minimum $S$ such that $v_{n+1}^{S^{\ast}}=0$ and Lemma 6, we now can state
@@
Appendix D Additional background
D.1 Extended literature review
@@
Recently, Alaa et al. (2023) provided predictive intervals for CATE meta-learners under the assumption of full knowledge of the propensity score. As an extension, Jonkers et al. (2024) proposed a Monte-Carlo sampling approach to receive less conservative intervals. Chen et al. (2024) provide prediction intervals for counterfactual outcomes. However, the proposed method requires access to additional interventional data and is thus not applicable to real-world applications on observational data. All methods are restricted to binary treatments.
Other works focus on prediction intervals for off-policy prediction (Taufiq et al., 2022; Zhang et al., 2023) and conformal sensitivity analysis (Yin et al., 2022), thus neglecting estimation errors arising from propensity or weight estimation or for randomized control trials (Kivaranovic et al., 2020; Lei & Candès, 2021). Wang et al. (2024) constructed intervals with treatment-conditional coverage of discrete treatments.
Aiming for group-conditional coverage, Wang et al. (2024) adapted CP to cluster randomized trials. Nevertheless, the method only applies to a finite number of treatments and thus is not applicable to continuous treatments.
@@
D.2 A note on challenges and difficulties in CP for causal effects of continuous treatments
@@
Existing works on conformal prediction for binary or (low-dimensional) discrete treatment are commonly based on (a) weighted conformal prediction (Tibshirani et al., 2019) or (b) conformal prediction local coverage guarantees (Lei & Wasserman, 2014). The first approach provides marginal coverage under a distribution shift through reweighting. It requires computing the weights based on the probability of treatment $A=a$. However, for continuous treatments, this is always zero. Although applicable to binary or low-dimensional discrete treatments (e.g., Lei & Candès, 2021), this weighting approach cannot be extended similarly to continuous treatments. Furthermore, the propensity of a continuous treatment given by the Dirac delta function $\delta_{a}$ would require us to restrict the calibration to data samples of the specific treatment, which are extremely rare or even might be missing. Therefore, the calibration step cannot be employed in our setting. The second approach provides treatment group-conditional coverage. Although again possible for binary or low-dimensional treatments, this approach does not apply to continuous treatments as no treatment groups can be defined. Instead, we propose a novel method for conformal predictions that circumvents the above problems and is carefully tailored to continuous treatments.
D.3 Causal effects of continuous treatments & kernel smoothing
@@
Causal inference becomes challenging with continuous treatments primarily due to the infinite number of potential outcomes per sample, from which only one outcome is observed. Continuous treatments thus result in causal effects that are generally represented by curves (called dose-response curves) (Kennedy.2018). This is unlike binary treatment, where the causal effects are represented by a single discrete value.
@@
For continuous treatments, the dose-response curves are typically assumed to fulfill some smoothness criterion (e.g., Patrick Schwab et al., 2020; Schweisthal et al., 2023). Hence, when estimating treatment effects, interpolation and kernel smoothing of the outcome function are commonly employed (e.g., Kennedy.2018; Nagalapatti et al., 2024).
@@
Underlying causal estimation with continuous treatments is the generalized propensity score (Imbens, 2000). It is defined as the conditional probability of receiving treatment $a^{\ast}$ given the covariates $X$ under the following regularity conditions: (i) For each $i$, $Y_{i}(a),x_{i},A_{i}$ are defined on a common probability space; (ii) $A_{i}$ is continuously distributed with respect to the Lebesgue measure; and (iii) $Y_{i}=Y_{i}(A_{i})$ is a well-defined random variable.
@@
Approximating the density $\delta_{a^{\ast}}(a)$ of the hard intervention $a^{\ast}$ through a Gaussian kernel follows directly from the definition of $\delta_{a^{\ast}}(a)$ as the limit of such kernel. This is also common in the literature (e.g., Kallus & Zhou, 2018). Importantly, we note that we do not directly approximate the potential outcome $Y(a^{\ast})$ (but only the propensity scores). Thus, we do not have a bias-variance trade-off of the estimated outcome. Due to the smoothness of the dose-response curve, it is now valid to employ observed samples within a treatment region of $a^{\ast}$ defined by $\sigma$ to construct the intervals. We note that the importance of the samples is weighted by the inverse distance of the sample to $a^{\ast}$ in treatment space. We give further intuition on the relationship between $\sigma$, the importance of observational samples, and the prediction interval width in the following.
D.4 Interpretation of optimal parameters
@@
In Theorem 2, we thus optimize over $\sigma>0$ and $c_{a}\in[\frac{1}{M},M]$ to obtain the $(1-\alpha)$-quantile of the distribution shift-calibrated non-conformity scores. The optimal parameter $\sigma^{\ast}$ represents a trade-off between the uncertainty in the prediction and the uncertainty in the interval construction: A small $\sigma^{\ast}$ resembles the propensity of the hard intervention best. Thus, with sufficient or even infinite data close to $a^{\ast}$, we could construct the narrowest CP interval. However, the smaller $\sigma$, the less data close to $a^{\ast}$ will be available to calculate the prediction interval in practice. As a result, many calibration data samples will be strongly perturbed during the calculation, which increases the uncertainty and, thus, the interval size.
Appendix E Experimentation details
E.1 Synthetic dataset generation

### v3

@@ 21      $S^{\ast}\leftarrow\frac{S_{\mathrm{up}}+S_{\mathrm{low}}}{2}$;
Below, we state a second algorithm that is applicable if the propensity score is known. In this case, a convex optimization solver can be used.
Computational complexity: The complexity of running our algorithms depends heavily on the employed optimization solver with complexity $\sigma_{s}(n_{c})$ (e.g., polynomial complexity for suitable [^7] A suitable solver refers to a solver designed to handle the constrained convex problem in our theorem convex solvers) and the size of the calibration dataset $n_{c}$. This might become costly for large-scale calibration datasets in practice. The outer algorithm has a time complexity of at most $O(\log(\frac{S_{up}-S_{low}}{\varepsilon})+1)$. Overall, our algorithm has a fixed complexity of $O(\log(\frac{S_{up}-S_{low}}{\varepsilon})+\sigma_{s}(n_{c}))$. The complexity of deriving intervals through MC-dropout or ensemble methods depends, however, on the number of MC samples or models, respectively. The latter thus scales with the precision of the intervals, which might be difficult to control. In our work, we use optimization as a tool to provide conformal prediction intervals. Future research should focus on developing more efficient optimization algorithms for this task.
@@
Appendix C Semi-synthetic experiments
To underline the effectiveness of our method, we perform additional experiments on the semi-synthetic TCGA dataset. The Cancer Genome Atlas (TCGA) dataset [56] consists of a comprehensive and diverse collection of gene expression data. The data was collected from patients with different cancer types. In our experiment, we consider the gene expression measurements of the 4,000 genes with the highest variability, which we employ as our features $X$. The study cohort consisted of a total of 9659 patients. We model a continuous treatment based on the sum of the 10 covariates with the highest variance and assign a treatment effect that is constant in the sum of the covariates.
As in the main paper, we construct CP intervals for different interventions and confidence levels $\alpha$. We state the empirical coverage of our method in Table 2 below. The prediction performance of the trained model on the hold-out test dataset is reported. We find that our method is highly effective.
Confidence level
Intervention $\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
$\Delta=0.5$ 0.9680 (0.0324) 0.8920 (0.0391) 0.8040 (0.0741)
$\Delta=1.0$ 0.9733 (0.0377) 0.9500 (0.0500) 0.7667 (0.0618)
$\Delta=1.5$ 0.9400 (0.0438) 0.8920 (0.0699) 0.8200 (0.0619)
Table 2: Coverage of the intervals from our CP method on the TCGA dataset. We report the mean followed by the standard deviation in apprentices.
We observe that our method consistently achieves the desired coverage. To evaluate the usefulness of the intervals, we also report the interval width in Table 3 below. The range of the outcomes was 2.0.
Confidence level
Intervention $\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
$\Delta=0.5$ 0.1003 (0.0331) 0.0843 (0.0252) 0.0683 (0.0169)
$\Delta=1.0$ 0.1017 (0.0420) 0.0877 (0.0349) 0.0642 (0.0200)
$\Delta=1.5$ 0.0930 (0.0272) 0.0822 (0.0260) 0.0674 (0.0180)
Table 3: Width of the intervals from our CP method on the TCGA dataset. We report the mean followed by the standard deviation in apprentices.
Appendix D Proofs
D.1 Proofs of the supporting lemmas
In the following, we prove Lemma 4.3 and Lemma 4.4 from our main paper.
Proof of Lemma 4.3
@@ D.1 Proofs of the supporting lemmas
Proof of Lemma 4.4
We first prove that the problem (PS) fulfills the linear independence constraint qualifications. For all $i=m+1,\ldots,n+1$, we denote the constraints of problem (PS) as
@@
As a result, the constraints are linearly independent. This property suffices for the KKT conditions to hold at any (local) optimum of (PS).To furthermore show that the KKT conditions are also sufficient for a global optimum, we show that (PS) is Type-I invex. An optimization problem with objective function $f(x)$ and constraints $g(x)<=0$ with $x\in\mathbb{R}^{n+1}$ is Type-I invex at $x_{0}$, if there exists $\nu(x,x_{0})\in\mathbb{R}^{n+1}$, such that
@@
[20]. In problem (PS), the gradients of the objective function and of each constraint $h_{i}$ for all $i,j=m+1,\ldots,n+1$ at $x_{0}$ are given by
@@
the definition of Type-I invexity holds for (PS). Thus, the KKT conditions are also sufficient for a global optimum.
D.2 Proof of Theorem 4.2
We prove Theorem 4.2 in three steps: (i) We show that function class :=$\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$ indeed satisfies Eq. (2) for the intervention $a^{\ast}=a+\Delta_{A}$ and rewrite Eq. (5) as a convex optimization problem. (ii) We retrieve the corresponding dual problem, derive a dual prediction set, and show the equality of the coverage guarantee of the dual and the primal prediction sets. (iii) We derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. For further theoretical background on the idea of the proof, we refer to Gibbs et al. [18].
@@
Observe that $\mathbb{E}[f(A,X)]=\theta$ for all $f\in\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. Therefore, Eq. (2) simplifies to
@@
Thus, $\mathcal{F}$ satisfies the propensity shift from Eq. (2) for the soft intervention $a^{\ast}=a+\Delta_{A}$. Following Lemma 4.1, we thus aim to find
@@
For a reference, see [18]. The Lagrangian of the primal problem states
@@
Recall from Lemma 4.1 that, for
@@
With the definition of $S^{\ast}$ as the maximum optimizer $\eta_{n+1}^{S}$ with $\eta_{n+1}^{S}<1-\alpha$ and Lemma D.1, it follows that
@@
D.3 Proof of Theorem 4.5
We follow the same outline as in the proof of Theorem 4.2 in Section D.2. In Lemma 4.3, we motivated the functional class of distribution shifts. Therefore, it is left to prove the coverage guarantee of $C_{n+1}$.
@@
Coverage guarantees. As in D.2, we fix some $\sigma>0$ and $c_{a}\in[\frac{1}{M},M]$ to obtain a specific $f(a,x):=\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp{\left(-\frac{(a_{i}-a^{\ast})^{2}}{2\sigma^{2}}\right)}}{\hat{\pi}(a_{i}\mid x_{i})}$. We further denote $\hat{g}(a,x)\in\mathcal{F}$ the optimal solution given by the optimal values $\hat{\sigma}$ and $\hat{c_{a}}$.
With the definition of $S^{\ast}$ as the minimum $S$ such that $v_{n+1}^{S^{\ast}}=0$ and Lemma D.2, we now can state
@@
Appendix E Additional background
E.1 Extended literature review
Uncertainty quantification for causal quantities
There exist various methods for uncertainty quantification of causal quantities. These are often based on Bayesian methods [e.g., 1, 21, 22, 24]. However, Bayesian methods require the specification of a prior distribution based on domain knowledge and are thus neither robust to model misspecification nor generalizable to model-agnostic machine learning models. Other methods only provide asymptotic guarantees [e.g., 25, 27]. The strength of conformal prediction, however, is to provide finite-sample uncertainty guarantees.
@@ E.1 Extended literature review
Recently, Alaa et al. [2] provided predictive intervals for CATE meta-learners under the assumption of full knowledge of the propensity score. As an extension, Jonkers et al. [27] proposed a Monte-Carlo sampling approach to receive less conservative intervals. Chen et al. [9] provide prediction intervals for counterfactual outcomes. However, the proposed method requires access to additional interventional data and is thus not applicable to real-world applications on observational data. All methods are restricted to binary treatments.
Other works focus on prediction intervals for off-policy prediction [49, 60] and conformal sensitivity analysis [58], thus neglecting estimation errors arising from propensity or weight estimation or for randomized control trials [30]. Wang et al. [54] constructed intervals with treatment-conditional coverage of discrete treatments. Aiming for group-conditional coverage, Wang et al. [54] adapted CP to cluster randomized trials. Nevertheless, the method only applies to a finite number of treatments and thus is not applicable to continuous treatments.
Lei and Candès [36] consider the estimated propensity by incorporating the estimation error as a TV-distance term in the coverage guarantees. However, for large TV-distances (close to 1), the proposed method can only construct intervals with a very limited coverage $\alpha\in(0,1-TV)$. Hence, the method is not suitable for applications in medical practice. Our method, however, can also construct intervals with high coverage guarantees for high estimation errors. An increased error will widen the prediction intervals instead of reducing the coverage guarantee. We consider our approach more suitable for medical practice, as one can visually inspect the intervals and decide on the suitability of the task at hand.
@@
Conformal prediction under covariate shift Multiple works on CP with marginal coverage under distribution shifts between training and test data have been introduced in the literature [e.g., 8, 11, 15, 16, 17, 18, 19, 36, 42, 50, 57]. Our setting also involves a distribution shift due to the intervention on the treatment but differs from the latter in that the true distribution shift is unknown.
Gibbs et al. [18] introduced an approach to derive CP intervals under unknown distribution shifts. It proves valid finite-sample prediction intervals for all distribution shifts in a finite-dimensional function class. However, the approach does not directly apply to causal inference settings. Nevertheless, our framework builds upon the work by Gibbs et al. [18] in that we re-frame the proposed approach to apply to the distribution shift induced through the intervention in causal effect estimation. In this setting, the distribution shift is captured by the shift of the propensity function. Adapting Gibbs et al. [18] to a causal inference setting requires carefully addressing the underlying challenges that come from computing CP intervals in a causal inference setting (e.g., propensity score estimation, hard/soft interventions), which we regard as our main novelty and which is of immediate practical relevance (e.g., in personalized medicine).
E.2 Discussion on the tightness of our CP intervals:
Our method builds upon the idea of CP to provide finite-sample coverage guarantees. Notably, standard CP does not provide intervals that are proven to be sharp. To our knowledge, there is no method that provides sharp/the tightest possible intervals for potential outcomes. Exploring the tightness of our CP intervals is an interesting and important direction for future research.
In practice, it is possible to observe the width, and thus the informativeness for decision-making, of the intervals. However, coverage guarantees cannot be observed. Therefore, we help the decision-maker by providing valid intervals. The decision-maker has to decide, case by case, if the returned intervals are beneficial for the problem at hand. Note that this aspect of the informativeness of intervals holds true for any uncertainty quantification method (including those to be proven to be sharp).
E.3 A note on challenges and difficulties in CP for causal effects of continuous treatments
@@
Existing works on conformal prediction for binary or (low-dimensional) discrete treatment are commonly based on (a) weighted conformal prediction [50] or (b) conformal prediction local coverage guarantees [35]. The first approach provides marginal coverage under a distribution shift through reweighting. It requires computing the weights based on the probability of treatment $A=a$. However, for continuous treatments, this is always zero. Although applicable to binary or low-dimensional discrete treatments [e.g., 36], this weighting approach cannot be extended similarly to continuous treatments. Furthermore, the propensity of a continuous treatment given by the Dirac delta function $\delta_{a}$ would require us to restrict the calibration to data samples of the specific treatment, which are extremely rare or even might be missing. Therefore, the calibration step cannot be employed in our setting. The second approach provides treatment group-conditional coverage. Although again possible for binary or low-dimensional treatments, this approach does not apply to continuous treatments as no treatment groups can be defined. Instead, we propose a novel method for conformal predictions that circumvents the above problems and is carefully tailored to continuous treatments.
E.4 Causal effects of continuous treatments & kernel smoothing
@@
Causal inference becomes challenging with continuous treatments primarily due to the infinite number of potential outcomes per sample, from which only one outcome is observed. Continuous treatments thus result in causal effects that are generally represented by curves (called dose-response curves) [29]. This is unlike binary treatment, where the causal effects are represented by a single discrete value.
@@
For continuous treatments, the dose-response curves are typically assumed to fulfill some smoothness criterion [e.g., 41, 46]. Hence, when estimating treatment effects, interpolation and kernel smoothing of the outcome function are commonly employed [e.g., 29, 38].
@@
Underlying causal estimation with continuous treatments is the generalized propensity score [23]. It is defined as the conditional probability of receiving treatment $a^{\ast}$ given the covariates $X$ under the following regularity conditions: (i) For each $i$, $Y_{i}(a),x_{i},A_{i}$ are defined on a common probability space; (ii) $A_{i}$ is continuously distributed with respect to the Lebesgue measure; and (iii) $Y_{i}=Y_{i}(A_{i})$ is a well-defined random variable.
@@
Approximating the density $\delta_{a^{\ast}}(a)$ of the hard intervention $a^{\ast}$ through a Gaussian kernel follows directly from the definition of $\delta_{a^{\ast}}(a)$ as the limit of such kernel. This is also common in the literature [e.g., 28]. Importantly, we note that we do not directly approximate the potential outcome $Y(a^{\ast})$ (but only the propensity scores). Thus, we do not have a bias-variance trade-off of the estimated outcome. Due to the smoothness of the dose-response curve, it is now valid to employ observed samples within a treatment region of $a^{\ast}$ defined by $\sigma$ to construct the intervals. We note that the importance of the samples is weighted by the inverse distance of the sample to $a^{\ast}$ in treatment space. We give further intuition on the relationship between $\sigma$, the importance of observational samples, and the prediction interval width in the following.
E.5 Interpretation of optimal parameters
@@
In Theorem 4.5, we thus optimize over $\sigma>0$ and $c_{a}\in[\frac{1}{M},M]$ to obtain the $(1-\alpha)$-quantile of the distribution shift-calibrated non-conformity scores. The optimal parameter $\sigma^{\ast}$ represents a trade-off between the uncertainty in the prediction and the uncertainty in the interval construction: A small $\sigma^{\ast}$ resembles the propensity of the hard intervention best. Thus, with sufficient or even infinite data close to $a^{\ast}$, we could construct the narrowest CP interval. However, the smaller $\sigma$, the less data close to $a^{\ast}$ will be available to calculate the prediction interval in practice. As a result, many calibration data samples will be strongly perturbed during the calculation, which increases the uncertainty and, thus, the interval size.
The parameter $c_{a}$ allows us to incorporate the estimation error in the propensity score. It represents a weighting of the propensity shift such that the $(1-\alpha)$-quantile of the non-conformity scores is increased with higher estimation error.
E.6 Interpretation of parameter $M$
Our optimization requires the specification of a parameter $M$, denoting a bound on the propensity estimation error. One can view the parameter $M$ as a type of sensitivity parameter. Therefore, we follow former work in causal inference and propose to incorporate domain knowledge to specify the parameter $M$ [e.g., 13, 48]. Another way of making use of the parameter $M$ is to observe how the intervals change for varying $M$. This indicates how much effect the propensity misspecification has on the prediction interval and can help in making reliable decisions. A third option to calibrate $M$ is to employ measures for epistemic uncertainty on top of the propensity estimate when there is no domain knowledge for specifying $M$.
E.7 A note on the stability of our method
In our experiments, one can observe some instability for certain privacy budgets and intervention combinations. This is likely due to the fact that the CP coverage guarantees are only marginal. Therefore, we might experience under- or over-coverage. However, we note that across all runs, our method, on average, achieves the desired coverage. These instabilities only occur in single settings. Furthermore, the variance in coverage of our method is much lower than the coverage variance of the baselines. A more in-depth analysis of the stability of the proposed method is left for future work.
Appendix F Experimentation details
F.1 Synthetic dataset generation